Displaying reports 421-440 of 1084.Go to page Start 18 19 20 21 22 23 24 25 26 End
Reports until 21:07, Friday 04 October 2024
Atoms General
Daniel Grass - posted 21:07, Friday 04 October 2024 (11889)
PICAS (2D MOT) Supports Machined

I designed and machined some supports for the PICAS 2D MOT. These will support the MOT and allow it to be attached with the bellows to the chamber holding the 3D MOT to trap and cool the Rb atoms. See attached photo and part files. The part is not meant to clamp the 2.75" flange but go around it.

Images attached to this report
Non-image files attached to this report
Lab Infrastructure General
Alex Ramirez - posted 13:23, Wednesday 02 October 2024 (11888)
File path for Office Signs Illustrator File

Here is where to find the Illustrator files for the McCuller Lab signage in Bridge Basement.

\Nextcloud\Lab\signs_and_nametags\Room_Names_Template.ai

There is also a Word document for printing out more name tags if you need another for your desk. 
 

Please contact me, Alex, before saving or modifying the adobe illustrator files or save a new one as your own. 

For additional signs to be made, please get in touch with me on mattermost

Lasers Laser Filter Cavity (GQuEST)
Daniel Grass - posted 18:24, Tuesday 01 October 2024 (11887)
Custom Based and Internally Tapped 10 Inch Flange Attached to Laser Filter Cavity Cubes

[Alex, Ian, Jeff, Daniel]

I one by one took out the 12 point hex screws that secured the custom base of the Laser Filter Cavity (LFC) Cube and inserted and moderately tightened nickel coated hex screws, with a smaller screw head heights, and washers. I then did this process for the 12 point hex screws that directly secured the Tapped 10" Flange and added washers. I then tightened the screws in a star pattern until I could no longer see the copper gasket. See attached photos.

To put the cube on the base, we lifted from the same 2 lifting points in this post. The cube naturally tilts such that the base goes down furthest when lowered again.

We lifted the 2nd LFC Cube in the same way as the first and repeated the same process. This time, I used torque wrenches to tighten the screws, incrementing by 4-5 Nm every time. The ultimate tightness needed to no longer see the copper gasket was 33 Nm. I recommend going by more than 4-5 Nm becuase it adds more opportunity to strip a screw. I nearly stripped one, but was able to tighten it to 30 Nm. We then put the cube on its base.

 

Images attached to this report
Lasers Laser Filter Cavity (GQuEST)
Daniel Grass - posted 20:23, Monday 30 September 2024 (11886)
Cube Tipping and End Cube Surgery

[Jeff, Ian, Lee, Daniel]

We needed to tip over an end cube onto a side so that we can replace the base flange with one with a breadboard pattern so that we can internally mount optics. We also need to do this to put on a custom base so the end cube can go on an optics table. To tilt the cube this, we attached eye bolts and shackels to two of the end cube lifting points and lifted the cube. We used 5 washers on each eyebolt so that the flared end made solid contact. After lifting the cube, we tipped it over onto its side by lifting the 5" long reducing flange (10" CFF to 8" CFF). It's now resting on a piece of foam that can support the cube and is long enough that the 14.5" diameter base isn't hitting the table.

We then unbolted the base, wrapped it in clean foil, and attached my custom base (with a copper gasket of course). We had some trouble getting the gasket in, but I eventaully got it by putting the gasket in the flange and some trial and error. We then added the 12 point silver coated 1.75" long 5/16-24 screws from duniway. There were three problems.

1. I forgot to add washers to the screws. Kurt J Lesker says its optional, but I think this is best practice and we have them.

2. The heads of the 12 point screws stick out just past the custom base.

3. The copper gasket was still visable after tightening to a 180 inch-lbs spec. I don't think this spec is correct and we should tighten more.

I ordered some zinc coated hex bolts from McMaster that should arrive tomorrow (I couldn't find silver coated hex bolts after looking at a few suppliers). We can solve all of these issues then.

Images attached to this report
Atoms General
Daniel Grass - posted 12:25, Monday 30 September 2024 (11885)
SM05RRV stuck in custom 3D MOT part and then removed

[Ian, Daniel]
 

While testing the bottom SM05 thread for the custom 3D MOT part, the SM05RRV got stuck. It would spin but not engage the correct threads to come out. A 0.5" end mill was able to destroy the ring for removal and keep the custom part in tact (in seems). I am going to retap all the threads.

Lab Infrastructure General
Daniel Grass - posted 12:23, Monday 30 September 2024 - last comment - 20:32, Monday 07 October 2024(11884)
B111A Oxygen Sensor

[Ian, Daniel, Sander]
 

Ian and I plugged in the East Bridge B111A O2 monitor, and it read 0.0 and then FLT then I unplugged it after a second when it alarmed.

 

Comments related to this report
Daniel Grass - 20:32, Monday 07 October 2024 (11892)

This seems to be fixed now.

Atoms General
Daniel Grass - posted 14:36, Friday 27 September 2024 (11882)
Cold Quanta PICAS a bit dusty out of box

I inspected the Cold Quanta PICAS (the Rb 2D MOT) out of the box. It was a bit dusty but mostly outside the vacuum area. I looks like my currently designed mounting structure will work.

 

Also of note is the damage sticker was triggered. 

Atoms General
Daniel Grass - posted 23:32, Thursday 26 September 2024 (11881)
Ideal Vac Vacuum Components Improperly Packaged likely require cleaning

3 of the items we recieved from Ideal Vac are not packaged to vacuum specifications and will likely require cleaning. These are the 6x6x6 in Cube, the 6x6x12 in Rectangular Prism, and the 6" Base. The 6" Base has metal shavings in the packaging. See attached photos.

Images attached to this report
Lab Infrastructure General
Daniel Grass - posted 18:48, Thursday 26 September 2024 (11880)
Items moved from B150 and B102 in B111

I moved various pieces of equipment from B150 and B102 into B111. I organized stuff as follows:

B111A North: RbQ Vacuum Equipment

B111A South: coated OFC cavity and Silicon parts

B111B: LFC parts and short Michelson parts

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:34, Wednesday 25 September 2024 (11879)
Bend Cubes Emptied and Inspected

[Jeff, Lee, Daniel]

We opened up the bend cubes from the holometer and removed their cylindrical bases, viton balls, and a few screws. We wrapped these parts in aluminum foil. We don't yet have a good system for labeling parts meant for UHV so we didn't label these parts (but it is quite obvious what is what, and I think these parts will never get used). 

The vacuum side of the base flange has 4 1/4-20 tapped holes. We could potentially use these flanges, but since we have to take off the bottom and add the custom base I made for mounting, it's probably not worth keeping these.

Both bend cubes have grime inside of them. They will have to be cleaned somewhat extensively.

Images attached to this report
Vacuum Systems General
Daniel Grass - posted 17:25, Wednesday 25 September 2024 (11878)
Stainless Steel Hex Keys cleaned to Class B UHV Standard

I cleaned the stainless steel hex keys to a Class B UHV Standard (normal cleaning but no bake out. LIGO does an air bakeout for 24 hours but I think this is unnecessary). Class B parts can come into contact with UHV parts but cannot be left in vacuum.

The smallest hex key in this set is 3/32". We need 5/64", 1/16", and 0.050" hex keys. McMaster has some Titanium hex keys. I didn't see any better options.

GQuEST General (GQuEST)
Alex Ramirez - posted 16:26, Wednesday 25 September 2024 (11877)
Fermilab and Dewar Update

During my visit to FNAL, we have been working toward the goal of prepping the Dewar for final shipment to Caltech.
When I arrived the Dewar was assembled and pumped down (without greasing any O-rings) to a leak rate of 10^-4 or 5. 


We then performed leak tests using helium and a helium detector to pinpoint a few locations for possible leaks.
Upon dismantling the entire device, I could locate a few particles of dust, debris, and fibers that may have caused these leaks and warked areas of concern with possible micro scratches or abrasions to the sealing surface of one of the inner cubes. We expect that on the second test, bringing the Dewar to vacuum to leak check for Thursday and sealing with vacuum grease, we will see a majority of these leaks disappear. Great care and meticulous cleaning were done in preparation for this cycle, so we will post the results on our leak rate tomorrow. 


Some other notes to take into consideration when I am to get back to Caltech:
- The copper cube for the "free space" optics has 2 notes.

     1. When manufacturing the plate for it, they were unable to meet the spec we initially required for the optics, thus each of the tapped holes is set to a tolerance of +-1 thousandth true position rather than 0.5 thousandths, this may or may not affect our alignments, but we will have to see when testing the dewar fully.

     2. These holes were made blind, they are through holes and will need to be blocked underneath with plugs or taped over with copper tape to block the light out. If this is a huge issue and the tape is not cutting it, we can also send it to the manufacturer for them to weld seal the holes. 

- All copper plates and components have been designed without helicoils as they would greatly affect the conduction inside the copper cubes. Thus we must take great caution when bolting anything together with the copper boxes to not strip any of the components!

- That being said, all thermal specs were done for bolts to have been tightened to 25 foot-lbs of force, we should buy some torque wrenches for the reassembly. 

- The copper is being passivated such that it is less likely to rust and is more durable over time, but this may need to be done again in the future (5-8 minute citric acid bath and 250-300 degree bakeout and cool down for 18 hours) 

- A few of the boxes being shipped will have not been opened before, so we will need to do a full inspection of the components inside them to alert the manufacturers if anything is wrong or damaged from shipment. 

- When the dewar is fully assembled it must not ever be tipped on its side as the G10 legs for the copper could be broken or damaged!

- Anytime we bring the dewar back to normal pressure (1 atm) we will want to infill with nitrogen, thus we will be adding a pressure relief valve to add a layer of safety when doing so and not overpressurizing the vessel. 
Things to check up on with Boris and the SNSPD group regarding components for the Dewar:

- Did we order the XYZ stage? We will also need to do a bit of testing with its internal fine-threaded bolts and their susceptibility to expanding in low-temperature environments (may need to have them made out of aluminum or steel instead of brass)

- Have we ordered the feed-throughs for the XYZ stage aligning?

- Suggested we order an extra set of the G10 feet (erik@precisioncryo.com) - they will have our order info for a bit and could replicate our parts (Fermi PO# PO711633)

Images attached to this report
Lab Infrastructure General
Daniel Grass - posted 01:04, Wednesday 25 September 2024 (11876)
B111 and B110 Cleaner

[Jeff, Daniel]
 

I moved parts out of the mobile clean room and wipes down the tables in side them. I also cleaned the outside of the end cubes to remove any anti sieze.
 

I then swiffered and vacuumed the floors of the entirety of B111 and B110. 
 

Jeff and I cleaned all of the B111A and B111B tables.
 

Jeff has cleaned all of the electronics racks.

Equipment Loans
Daniel Grass - posted 14:48, Monday 23 September 2024 - last comment - 12:56, Friday 11 October 2024(11875)
Silver Plated 12 Point Bolts Loaned to Paco

I loaned a bag of Silver Plated 12 Point Bolts (1/4-28 x 7/8") with 1/4" washers, 25/pk to Paco. Duniway Part Number SBX-28-087. Paco said he would return the bag on Thursday, 9/26.

Comments related to this report
Daniel Grass - 12:56, Friday 11 October 2024 (11898)

A new bag of this equipment has been traded to us.

Computing General
Jeffrey Wack - posted 17:14, Sunday 22 September 2024 (11874)
Installing Debian 11 (Bullseye) on CyMAC front-end machine

I have installed Debian 11 on the server which is currently the only machine in the mobile server rack in B111D. Here is a guide to the steps I took.

1) create a bootable usb drive. This tutorial is very detailed. Here is the downloads page for Debian 11 (bullseye). Use the netinst CD image.

2) Plug into the machine: keyboard, monitor, ethernet connection to network, and bootable usb drive (also called installation media)

3) boot up the machine and press F11 when the option appears to enter a boot menu

4) select your usb drive from the list and then follow the installation wizard

5) when asked to select a network I was given the options 'eno1' and 'eno2', only eno2 worked (I do not know what these two options refer to)

6) I selected the hostname 'babbage' and username 'cymac'. The root and user passwords are on the labsecrets wiki page.

7) When asked about disk partitioning, I did a 'guided' partition of the entire disk, and did not modify any of the defaults.

8) When asked what software you'd like, unselect "Debian Desktop Environment" and "GNOME" and select "MATE". Leave "Standard tools" selected. Installing GNOME led to graphics related crashes down the line.

That's all! There will be a future post about configuring Bios settings and installing CyMAC software.

 

Environmental Dust/Particulates
Daniel Grass - posted 18:15, Thursday 19 September 2024 - last comment - 14:48, Monday 13 January 2025(11873)
Particle Count Measurements in B102, B111, B111A, and B111B

I took some particle counter measurments in B102, B111, B111A, and B111B. In all the clean rooms, the particle count was 0. These clean rooms were the B102 cleanroom with the optical cavities, the B111B mobile cleanrooms, the Southern most B111A long table, and the Northern most B111B short table.

I also took some 10 second particle counts outside the clean rooms. Data is 0.3 um count/0.5 um count/1.0 um count)

B102: 1268/72/28 (there were lots of people and activity during this measurment)

B111 (entryway): 495/46/28

B111B: 177/51/39

B111A: 95/17/13

These seem like very good measurments.

The particle counter is on the optics tables in B111B.

Comments related to this report
Daniel Grass - 19:58, Sunday 12 January 2025 (12070)

It should probably read "Northern most B111A short table" considering there is no short table in B111B.

Daniel Grass - 14:48, Monday 13 January 2025 (12072)

Data extrapolated to a full minute for easy comparisons:

B102: 7608/432/168 (there were lots of people and activity during this measurment)

B111 (entryway): 2970/276/168

B111B: 1062/306/234

B111A: 564/102/78

Lasers Laser Filter Cavity (GQuEST)
Daniel Grass - posted 18:10, Thursday 19 September 2024 (11872)
Holometer End Cubes designated for LFC Moved to B111B

[Jeff, Ian, Torrey, Alex, Daniel]

We turned off the lasers in B102 for safety purposes (we'd be opening and closing the curtains frequently) and moved both Holometer End Cubes designated for the LFC into B111B onto a 2' tall table.

The next steps are cleaning B111B (with special attention to the mobile clean rooms) and replacing the bottom flange with one with a breadboard pattern.

Images attached to this report
Lasers General
Torrey Cullen - posted 17:02, Thursday 19 September 2024 - last comment - 17:03, Thursday 19 September 2024(11870)
B102 is Laser Safe

The 1550 seed laser and amplifier in B102 are now off. The sign on the outside of the room is also off.

Comments related to this report
Daniel Grass - 17:03, Thursday 19 September 2024 (11871)

The 780 nm RbQ laser has remained off.

Equipment General
Torrey Cullen - posted 12:58, Thursday 19 September 2024 (11869)
Particle Counter

The particle counter calibration certificate expires October 15th. The company reached out for a recalibration service. We've declined it at this time. 
Point of contact: Sophia Correia (srcorreia@beckman.com)

Displaying reports 421-440 of 1084.Go to page Start 18 19 20 21 22 23 24 25 26 End